I need to find a linear equation for the forest area in the sudan. In 1990 the forest area was 763814 sq.Km. In 2005 the forest area was 675458 sq km. What linear equation would show the decrease in forest area. Furthermore, what will the forest area be in 2010?
Answer by stanbon(48568)

I need to find a linear equation for the forest area in the sudan.
In 1990 the forest area was 763814 sq.Km.
In 2005 the forest area was 675458 sq km.
What linear equation would show the decrease in forest area.
Furthermore, what will the forest area be in 2010?
--------------------------
Let "x" correspond to 1990.
-----
You have two points so find the slope.
--------------
slope (763814-675458)/(0-15) = -5890.4
--------
The intercept is 763814 because that corresponds to x=0.
---------------------------------
Equation:
Forest area in year x = -5690.4(x) +763814
------------
what will the forest area be in 2010?
FA(20) = -5690.4*20 + 763814
FA(20) = 646,006 acres (forest area in 2010)
====================================================
Cheers,
Stan H.